a wire is bent in the form of a quadrant of a circle of diameter 1.4m. find the length of the wire, if the quadrant is bounded by the radii of the circle.
Answer #1
The radii are 0.7 each and the length of the arc is 1/4 of the circumference. pi*D/4=pi*1.4/4=1.1 1.1+.7+.7=2.2 m Ed
answered by: Jkim